OpenStudy (anonymous):

A scale on a map says that 1 inch represents 60 miles. If the distance between two points on the map is 2 1/2 inches, how many miles are between the two points? 150 120 24 180

OpenStudy (anonymous):

@ganeshie8

Elsa213 (elsa213):

The answer is 150 /.<

Elsa213 (elsa213):

m/r = 1 in / 60 miles 1/60 = 2.5/r r = 2.5(60/1) c: http://www.algebra.com/algebra/homework/word/travel/Travel_Word_Problems.faq.question.817829.html